# PVChat: Personalized Video Chat with One-Shot Learning

### v3

@@ Abstract
Video large language models (ViLLMs) excel in general video understanding, e.g., recognizing activities like talking and eating, but struggle with identity-aware comprehension, such as “Wilson is receiving chemotherapy” or “Tom is discussing with Sarah”, limiting their applicability in smart healthcare and smart home environments. To address this limitation, we propose a one-shot learning framework PVChat, the first personalized ViLLM that enables subject-aware question answering (QA) from a single video for each subject. Our approach optimizes a Mixture-of-Heads (MoH) enhanced ViLLM on a synthetically augmented video-QA dataset, leveraging a progressive image-to-video learning strategy. Specifically, we introduce an automated augmentation pipeline that synthesizes identity-preserving positive samples and retrieves hard negatives from existing video corpora, generating a diverse training dataset with four QA types: existence, appearance, action, and location inquiries. To enhance subject-specific learning, we propose a ReLU Routing MoH attention mechanism, alongside two novel objectives: (1) Smooth Proximity Regularization for progressive learning through exponential distance scaling and (2) Head Activation Enhancement for balanced attention routing. Finally, we adopt a two-stage training strategy, transitioning from image pre-training to video fine-tuning, enabling a gradual learning process from static attributes to dynamic representations. We evaluate PVChat on diverse datasets covering medical scenarios, TV series, anime, and real-world footage, demonstrating its superiority in personalized feature understanding after learning from a single video, compared to state-of-the-art ViLLMs.
@@ 1 Introduction
Recent advancements in Video Large Language Models (ViLLMs) have showcased impressive capabilities across a wide range of tasks, e.g., video captioning, temporal action localization, and video question answering [5, 50, 31, 49]. These models, trained on vast datasets, exhibit broad knowledge across multiple domains, such as medical assistance [17] and autonomous driving [55], significantly extending the functional scope of ViLLMs.
@@
  • We develop a systematic video augmentation and QA generation pipeline from a single reference video. The key is to exploit off-the-shelf toolboxes to generate identity-preserving positives and synthesize hard negatives for confusingly similar faces. To support the research, we construct a diverse dataset containing $6$ individual scenarios, $304$ original videos, $2{,}304$ extended generated videos with over $30{,}000$ QA pairs, which will be publicly released to facilitate future research.
@@ 3.2 Data Collection
Through this comprehensive pipeline, each input natural video is transformed into $81$ different videos accompanied by $1{,}455$ question-answer pairs, significantly expanding our training dataset while maintaining subject identity consistency and guaranteeing the model’s robustness.
@@ 3.3 ReLU Routing Mixture-of-Heads Attention
[…]
@@ 3.3.2 Controlling Sparsity Strategy
However, $L_{Reg}$ just focuses on increasing the sparsity, forcing the model to be more sparsely activated, which usually makes all output prone to 00. Considering this, we design a Head Activation Enhancement (HAE) Loss to activate more heads to avoid some experts always being asleep:
@@ 4.1 Settings
[…]
Datasets. To establish the dataset, we collect six different scenarios, including (Friends(6), Good Doctor(5), Ne Zha(2), doctor(3), patient(3), Big Bang(6)) $25$ characters (including doctor, patient, cartoon character, engineer, professor, etc.), and collect $304$ original videos and $2{,}304$ extended videos with over $30{,}000$ QA pairs. $300$ prompt descriptions for different scenarios are designed for different ages and genders. The structure and distribution of our dataset can be found in Fig. 6. For each character, we randomly select one video for the training set as our refer video and the other for the evaluation set as query video. We show more details and examples in our supplementary.
@@ 5 Conclusion
This paper introduces PVChat, the first personalized ViLLM that enables personalized subject learning from a single video. We collect comprehensive datasets over $2{,}300$ videos, propose ReMoH for specific knowledge learning, and employ a two-stage training strategy from image pre-training to video fine-tuning, with SPR and HAE losses. Evaluations show the superiority of our model in personalized video QA, making it potential for smart healthcare and home scenarios.
@@
[…]

### v4

@@ Abstract
Video large language models (ViLLMs) excel in general video understanding, e.g., recognizing activities like talking and eating, but struggle with identity-aware comprehension, such as “Wilson is receiving chemotherapy” or “Tom is discussing with Sarah”, limiting their applicability in smart healthcare and smart home environments. To address this limitation, we propose a one-shot learning framework PVChat, the first personalized ViLLM that enables subject-aware question answering (QA) from a single video for each subject. Our approach optimizes a Mixture-of-Heads (MoH) enhanced ViLLM on a synthetically augmented video-QA dataset, leveraging a progressive image-to-video learning strategy. Specifically, we introduce an automated augmentation pipeline that synthesizes identity-preserving positive samples and retrieves hard negatives from existing video corpora, generating a diverse training dataset with four QA types: existence, appearance, action, and location inquiries. To enhance subject-specific learning, we propose a ReLU Routing MoH attention mechanism, alongside two novel objectives: (1) Smooth Proximity Regularization for progressive learning through exponential distance scaling and (2) Head Activation Enhancement for balanced attention routing. Finally, we adopt a two-stage training strategy, transitioning from image pre-training to video fine-tuning, enabling a gradual learning process from static attributes to dynamic representations. We evaluate PVChat on diverse datasets covering medical scenarios, TV series, anime, and real-world footage, demonstrating its superiority in personalized feature understanding after learning from a single video, compared to state-of-the-art ViLLMs. Code is at PVChat.
@@ 1 Introduction
Recent advancements in Video Large Language Models (ViLLMs) have showcased impressive capabilities across a wide range of tasks, e.g., video captioning, temporal action localization, and video question answering [5, 50, 31, 49]. These models, trained on vast datasets, exhibit broad knowledge across multiple domains, such as medical assistance [17] and autonomous driving [55, 56], significantly extending the functional scope of ViLLMs.
@@
  • We develop a systematic video augmentation and QA generation pipeline from a single reference video. The key is to exploit off-the-shelf toolboxes to generate identity-preserving positives and synthesize hard negatives for confusingly similar faces. To support the research, we construct a diverse dataset containing $6$ individual scenarios, $304$ original videos, 2,3042{,}304 extended generated videos with over 30,00030{,}000 QA pairs, which will be publicly released to facilitate future research.
@@ 3.2 Data Collection
Through this comprehensive pipeline, each input natural video is transformed into $81$ different videos accompanied by 1,4551{,}455 question-answer pairs, significantly expanding our training dataset while maintaining subject identity consistency and guaranteeing the model’s robustness.
@@ 3.3 ReLU Routing Mixture-of-Heads Attention
 $\begin{array}[]{l}\text{ Attention }(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\operatorname{Softmax}\left(\frac{\boldsymbol{Q}\boldsymbol{K}^{\top}}{\sqrt{d_{k}}}\right)\boldsymbol{V},\\ \text{where}\>\boldsymbol{Q}=\boldsymbol{X}_{1}\boldsymbol{W}_{Q},\boldsymbol{K}=\boldsymbol{X}_{2}\boldsymbol{W}_{K},\boldsymbol{V}=\boldsymbol{X}_{2}\boldsymbol{W}_{V}.\end{array}$ (1)
@@ 3.3.2 Controlling Sparsity Strategy
However, $L_{Reg}$ just focuses on increasing the sparsity, forcing the model to be more sparsely activated, which usually makes all output prone to 0. Considering this, we design a Head Activation Enhancement (HAE) Loss to activate more heads to avoid some experts always being asleep:
@@ 4.1 Settings
Implementation Details. Our LLM backbone is based on the Mistral-7B-Instruct-v0.3 [11]. All models are trained on one NVIDIA L20 GPU. The first stage of training takes one epoch, and the second stage takes seven epochs. We set the batch size to $2$, the learning rates of token embedding, ReMoH, LLM Lora to $\times 10^{-4}, 1×1$$\times 10^{-5}$, and $1\times 10^{-5}$, training time needs $3h$, video resolution is $1080\times 1920$.For visual encoder each video will evenly sample 8 frames.
Datasets. To establish the dataset, we collect six different scenarios, including (Friends(6), Good Doctor(5), Ne Zha(2), doctor(3), patient(3), Big Bang(6)) $25$ characters (including doctor, patient, cartoon character, engineer, professor, etc.), and collect $304$ original videos and 2,3042{,}304 extended videos with over 30,00030{,}000 QA pairs. $300$ prompt descriptions for different scenarios are designed for different ages and genders. The structure and distribution of our dataset can be found in Fig. 6. For each character, we randomly select one video for the training set as our refer video and the other for the evaluation set as query video. We show more details and examples in our supplementary.
@@ 5 Conclusion
This paper introduces PVChat, the first personalized ViLLM that enables personalized subject learning from a single video. We collect comprehensive datasets over 2,3002{,}300 videos, propose ReMoH for specific knowledge learning, and employ a two-stage training strategy from image pre-training to video fine-tuning, with SPR and HAE losses. Evaluations show the superiority of our model in personalized video QA, making it potential for smart healthcare and home scenarios.
@@
ReMoH Within the ReMoH architecture, we implement cross-attention mechanisms at alternating layers to facilitate interaction between video embeddings (extracted by the visual encoder) and query embeddings. Our approach adaptively scales with the number of subjects: for single-person training, we incorporate an additional $16\times\text{token\_length}\times 768$ dimensional feature space; for two-person training, this expands to $32\times\text{token\_length}\times 768$ and for three-person training, it further increases to $48\times\text{token\_length}\times 768$.